OpenStudy (anonymous):

Which set of measurements would NOT form a triangle? a) 7, 12, 12 b) 7, 12, 5 c) 7, 12, 17 d) 7, 12, 7

OpenStudy (anonymous):

For the measurements to be good, two sides added together have to be greater than the other side. The only one that doesn't work here would be B, because 7+5=12 abd 12 isn't greater than 12. So B would not form a triangle.

OpenStudy (anonymous):

B wouldn't because the two sides must be greater when added up then the biggest angle

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

so its c ?

OpenStudy (anonymous):

No, the greatest angle measure has to be less than the other two added up. so C cant be it because 7+12>17 so it could form a triangle

OpenStudy (anonymous):

but 7+12= 19 > 17

OpenStudy (anonymous):

Exactly, when the sum is greater, which in this case 19>17 it means that it can form a triangle. This question is asking which one cannot form a triangle.

OpenStudy (anonymous):

omg, i didnt see the not lol

OpenStudy (anonymous):

Its fine good work, Im glad that you know have an understanding of the material. Have a good day!

OpenStudy (anonymous):

none -.-

OpenStudy (anonymous):

B wouldnt because 7+5=12, the sum needs to be greater not equal.